# Infinite-Dimensional Feature Interaction

### v2

@@ 1 Introduction
In the past decade, deep neural network architecture design has experienced several major paradigm shifts regarding the feature representation learning. As shown in Fig. 1(a), the early stage of neural network design is dominant by flat stream architectures in the form of weight-feature interaction (e.g., $W\textbf{x}$), like multi-layer perceptron (MLP), convolution neural networks (CNN), ResNet, etc. These models usually adopt linear superposition (e.g., $W_{i}\textbf{x}\oplus W_{j}\textbf{x}$) in the feature representation space. Therefore, the feature representation space scaling is limited to increase model channel width and depth [16, 12]. Nevertheless, this scaling approach has witnessed model development from the very small-scale MLPs or LeNet[17] to the recent huge ConvNext V2 [37]. With the ultra-scaled parameter amounts, computing complexity, and model size, the return of investment on model performance by further scaling feature dimensions has largely plateaued [9].
@@
Despite the plateau in feature representation space, recent sporadic architecture design works [26] shed light on another potential dimension of scaling: feature interaction space. Specifically, as shown in Fig. 1(b), these neural network designs generally demonstrate feature-feature interaction (e.g., $W_{i}\textbf{x}\otimes W_{j}\textbf{x}$). As a mathematical example, the self-attention mechanism in Transformers [34] can be formulated as $\textbf{x}^{L+1}=f_{k}(\textbf{x}^{L})*f_{q}(\textbf{x}^{L})*f_{v}(\textbf{x}^{L})$, which is also element-wise multiplication between processed input feature themselves. Characterized by element-wise interaction, these designs offer complementary feature correlation capabilities in addition to simple linear superposition. Such feature interactions have become the essential mechanisms of mainstream state-of-the-art neural architectures. For example, it’s implemented in SENet with squeeze and excitation [15], non-local network with transposed multiplication [35], vision transformers with self-attention [34, 3, 10], gated aggregation [19, 31, 42], and quadratic neurons [40, 11, 41].
@@
Although these models greatly improve the performance of state-of-the-arts, as mentioned above, these works provide diversified explanations that neglect the underlying shared design of element-wise feature multiplication operation [26], and thus may fail to reveal the fundamental source of improvement. To provide both explainability and quantifiability, in this paper, we propose a unified theoretical perspective to rethink the feature interaction scaling, i.e., the dimensionality of feature interaction space. For example, as shown in Fig 1(b), by employing the $\otimes$ multiplication, an implicit interactive quadratic space $\mathcal{Q}=\text{span}\left(x_{1}^{2},x_{1}x_{2},\cdots,x_{n-1}x_{n},x_{n}^{2}\right)$ with degrees of freedom ${n(n+1)}/{2}$ is constructed from the original representative vector space $\mathcal{V}=\text{span}\left(x_{1},x_{2},\cdots,x_{n}\right)$ with degrees of freedom $n$ [26]. Such space dimensionality scaling is the key to improving feature representation quality and end-task, as we will show later.
@@
From the unified perspective, a new opportunity emerges in neural architecture design, that is to scale to infinite-dimensional feature interactions than former methodologies (e.g., from $n(n+1)/2$ to $\lim_{k\to\infty}\frac{(n+k-1)!}{(n-1)!k!}$). However, scaling feature interaction space dimensionality from architectural enhancements (e.g., quadratic, self-attention, and recursive gates) comes with linear scaling cost w.r.t. order $k$, which hinders the infinite dimensionality increase [19, 10]. Thus, there is an open question:
@@
We then propose InfiNet, a novel family of neural networks that generate high-quality feature representations. Specifically, we introduce the Radial Basis Function (RBF) kernel [29] as a replacement of the common $\oplus$ or $\otimes$ operations. With the infinite series expansion in RBF kernel, it enables a theoretical provable dimensionality approximation $\text{span}_{j\in\mathbb{N},\sum_{k}{n_{k}}=j}\{\frac{x_{1}^{n_{1}}\cdots x_{k}^{n_{k}}}{\sqrt{n_{1}!\cdots n_{k}!}}\}$ while with as low-overhead as exponential operations. In this way, InfiNet enables efficient infinite-dimensional feature interaction space scaling upon a finite set of branches in the model architecture, thus achieving better complexity performance tradeoffs than prior state-of-the-arts.
@@ 2 Related Work
Kernel Method in Neural Networks. The fundamentals of kernel methods are well-studied in traditional ML domains like support vector machines [32] but they are less used in neural architectures. Early extensions to deep learning through the kernelized perceptron [6] have improved the performance but mainly in shallow neural networks. Recent advancements include the development of Convolutional Kernel Networks (CKN) [27], which merge CNNs’ robust feature learning with kernel stability. This approach offers a theoretical foundation for deep learning’s application in structured data. Additionally, the introduction of Kervolutional Neural Networks [5] replaces traditional convolution in CNNs with kernel-based operations to enhance feature extraction without excessive computational costs. The combination of Gaussian processes with neural networks to create Deep Kernel Learning [36] adjusts model complexity based on data while maintaining Bayesian inference.
@@ 3 From Feature Representation Space to Interaction Space
We start with considering the transformation of feature representation of a normal isotropic 2-layer perceptron block. Given an input $\mathbf{x}=(x_{1},x_{2},\cdots,x_{n})$ with n-dimension. We denote the first layer transformation as $\mathbf{z}=g(\mathbf{x})=\sigma(W_{1}\mathbf{x})$ (we omit the bias term for simplicity), and the second layer transformation as $h(\mathbf{z})=W_{2}\mathbf{z}$. Therefore the whole block is a mapping $h\circ g:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ from feature space $\mathbb{R}^{n}$ to a medium feature representation space $\mathbb{R}^{n}$ and eventually to a output space $\mathbb{R}^{n}$. Expanding the width of the neural networks, for example with a coefficient 2, is going to expand the feature representation space to $\mathbb{R}^{2n}$. The core idea of this dimensional expansion is that implicit associations in feature in low dimensions will be expressed explicitly when projected into a high-dimensional space. Model architecture (i.e. convolution, multi-branch [33]) is an additive superposition of pixels, essentially no different from the perceptron in terms of representation space.
@@ 3.2 Dimension of Feature Interaction Space
For example, we consider the Self-Attention in the Transformers [34]. The Self-Attention contains two element-wise multiplication, so it explores a 3-order feature interaction space. This is due to the fact that: (1) in first stage, in the query-key dot-product attention map computation $Att(\mathbf{x})=Q\cdot K^{T}=W_{Q}\mathbf{x}\cdot\mathbf{x}^{T}W_{K}^{T}$ explores the feature interaction space $\mathcal{A}=\text{span}\left(x_{1}^{2},x_{1}x_{2},\cdots,x_{n-1}x_{n},x_{n}^{2}\right)$. (2) In the second stage, the multiplication between the attention map and the value $y=\sigma(Att(\textbf{x}))\cdot V=\sigma(Att(\textbf{x}))\cdot W_{v}\mathbf{x}$ explores the feature interaction space $\mathcal{S}=\mathcal{A}\otimes\mathbb{R}^{n}=\text{span}\left(x_{1}^{3},x_{1}^{2}x_{2},x_{1}x_{2}x_{3},\cdots,x_{n-1}x_{n}^{2},x_{n}^{3}\right)$, which has $(n+2)(n+1)n/6$ dimensions in line with Eq.(5). Thereby we explain, from the perspective of feature interaction space, why the transformer family of models has, so far, generally outperformed recurrent neural networks and convolutional neural networks in various domains.
@@ 5.1 InfiBlock: Infinite-Dimensional Spatial Feature Interaction
where $K_{rbf}(\vec{Z}^{l}_{a},\vec{Z}^{l}_{b})=exp(-\frac{\|\vec{Z}_{a}^{l}-\vec{Z}_{b}^{l}\|_{2}^{2}}{2})$, is an RBF kernel with out any hyper-parameter, and $\|\cdot\|_{2}^{2}$ is squared L-2 norm. $\text{LayerNorm}(\cdot)$ is layer normalization [2] and $\text{MLP}(\cdot)$ is activated by GELU [13].
@@ 6.1 ImageNet Classification
Results. We present our ImageNet experiment results and comparison with baselines in Table 1. Our models achieve competitive performance with state-of-the-art. It is worth noting that InfiNet has about 20% fewer FLOPs than the baseline models with similar parameter scales, but still achieves great performance. It demonstrates the effectiveness of our proposed generation of features in infinite-dimensional interaction spaces. Experiments on isotropic models can be found in Table 3(a).
@@ A.2 More Visualization Comparison
A.3 Limitations
Although we explored polynomial kernels and RBF kernels, there are still a wide variety of kernel methods that we did not explore, including a learnable kernel. In addition, to avoid additional hyperparameter tuning, we fixed the $\sigma$ parameter in the RBF kernel to 1. This may have deprived us of the possibility of exploring the optimal InfiNet, but due to the high cost of training the model, we will leave the impact of this hyperparameter on the InfiNet as a follow-up work.
A.4 Broader Social Impact

### v4

@@ 1 Introduction
In the past decade, deep neural network architecture design has experienced several major paradigm shifts regarding the feature representation learning. As shown in Fig. 1(a), the early stage of neural network design is dominant by flat stream architectures in the form of weight-feature interaction (e.g., $W\textbf{x}$), like multi-layer perceptron (MLP), convolution neural networks (CNN), ResNet, etc. These models usually adopt linear superposition (e.g., $W_{i}\textbf{x}\oplus W_{j}\textbf{x}$) [^1] The symbol $\oplus$ denotes the Direct Sum of vector spaces. This corresponds to structures such as channel expansion/bottleneck in commonly used neural networks. While the $\otimes$ is elementwise multiplication. in the feature representation space. Therefore, the feature representation space scaling is limited to increase model channel width and depth [19, 14]. Nevertheless, this scaling approach has witnessed model development from the very small-scale MLPs or LeNet[20] to the recent huge ConvNext V2 [41]. With the ultra-scaled parameter amounts, computing complexity, and model size, the return of investment on model performance by further scaling feature dimensions has largely plateaued [10].
@@
Despite the plateau in feature representation space, recent sporadic architecture design works [29] shed light on another potential dimension of scaling: feature interaction space. Specifically, as shown in Fig. 1(b), these neural network designs generally demonstrate feature-feature interaction (e.g., $W_{i}\textbf{x}\otimes W_{j}\textbf{x}$). As a mathematical example, the self-attention mechanism in Transformers [38] can be formulated as $\textbf{x}^{L+1}=f_{k}(\textbf{x}^{L})\otimes f_{q}(\textbf{x}^{L})\otimes f_{v}(\textbf{x}^{L})$, which is also element-wise multiplication between processed input feature themselves. Characterized by element-wise interaction, these designs offer complementary feature correlation capabilities in addition to simple linear superposition. Such feature interactions have become the essential mechanisms of mainstream state-of-the-art neural architectures. For example, it’s implemented in SENet with squeeze and excitation [17], non-local network with transposed multiplication [39], vision transformers with self-attention [38, 4, 11], gated aggregation [22, 35, 46], and quadratic neurons [44, 12, 45].
@@
Although these models greatly improve the performance of state-of-the-arts, as mentioned above, these works provide diversified explanations that neglect the underlying shared design of element-wise feature multiplication operation [29], and thus may fail to reveal the fundamental source of improvement. To provide both explainability and quantifiability, in this paper, we propose a unified theoretical perspective to rethink the feature interaction scaling, i.e., the dimensionality of feature interaction space. For example, as shown in Fig 1(b), by employing the $\otimes$, element-wise multiplication, an implicit interactive quadratic space $\mathcal{Q}=\text{span}\left(x_{1}^{2},x_{1}x_{2},\cdots,x_{n-1}x_{n},x_{n}^{2}\right)$ with degrees of freedom ${n(n+1)}/{2}$ is constructed from the original representative vector space $\mathcal{V}=\text{span}\left(x_{1},x_{2},\cdots,x_{n}\right)$ with degrees of freedom $n$ [29]. Such space dimensionality scaling is the key to improving feature representation quality and end-task, as we will show later.
@@
From the unified feature dimensionality perspective, a new opportunity emerges in neural architecture design, that is to scale to infinite-dimensional feature interactions than former methodologies (e.g., from $n(n+1)/2$ to $\lim_{k\to\infty}\frac{(n+k-1)!}{(n-1)!k!}$). However, scaling feature interaction space dimensionality from architectural enhancements (e.g., quadratic, self-attention, and recursive gates) comes with linear scaling cost w.r.t. interaction order $k$, which hinders the infinite dimensionality increase [22, 11]. Thus, there is an open question:
@@
We then propose InfiNet, a novel family of neural networks that generate high-quality feature representations. Specifically, we introduce the Radial Basis Function (RBF) kernel [32] as a replacement of the common $\oplus$ or $\otimes$ operations. With the infinite series expansion in RBF kernel, it enables a theoretical provable dimensionality approximation $\text{span}_{j\in\mathbb{N},\sum_{k}{n_{k}}=j}\{\frac{x_{1}^{n_{1}}\cdots x_{k}^{n_{k}}}{\sqrt{n_{1}!\cdots n_{k}!}}\}$ while with as low-overhead as evaluating an exponential function. In this way, InfiNet enables efficient infinite-dimensional feature interaction space scaling upon a finite set of branches in the model architecture, thus achieving better complexity performance tradeoffs than prior state-of-the-art.
@@ 2 Related Work
Kernel Methods in Neural Networks. The fundamentals of kernel methods are well-studied in traditional ML domains like support vector machines [36] but they are less used in neural architectures. Early extensions to deep learning through the kernelized perceptron [7] have improved the performance but mainly in shallow neural networks. Recent advancements include the development of Convolutional Kernel Networks (CKN) [30], which merge CNNs’ robust feature learning with kernel stability. This approach offers a theoretical foundation for deep learning’s application in structured data. Additionally, the introduction of Kervolutional Neural Networks [6] replaces traditional convolution in CNNs with kernel-based operations to enhance feature extraction without excessive computational costs. The combination of Gaussian processes with neural networks to create Deep Kernel Learning [40] adjusts model complexity based on data while maintaining Bayesian inference. The neural tangent kernel (NTK) [18] framework establishes a direct connection between infinitely wide neural networks at initialization and kernel methods. Specifically, NTK shows that as the width grows, the network’s training dynamics can be described by a kernel function, linking the neural network’s behavior to that of kernel methods. Random features [34] provide an efficient approximation to the feature mappings used in kernel methods. By random projections, one can approximate the inner product defined by a kernel function, making it feasible to apply kernel methods [13, 3].
@@ 3 From Feature Representation Space to Interaction Space
We start with considering the transformation of feature representation of a normal shape-preserving 2-layer perceptron block. Given an input $\mathbf{x}=(x_{1},x_{2},\cdots,x_{n})$ with n-dimension. We denote the first layer transformation as $\mathbf{z}=g(\mathbf{x})=\sigma(W_{1}\mathbf{x})$ (we omit the bias term for simplicity), and the second layer transformation as $h(\mathbf{z})=W_{2}\mathbf{z}$. Therefore the whole block is a mapping $h\circ g:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ from feature space $\mathbb{R}^{n}$ to a middle feature representation space $\mathbb{R}^{n}$ and eventually to a output space $\mathbb{R}^{n}$. Expanding the width of the neural networks, for example with a coefficient 2, is going to expand the feature representation space to $\mathbb{R}^{2n}$. The core idea of this dimensional expansion is that implicit associations in features in low dimensions will be expressed explicitly when projected into a high-dimensional space. Model architecture (i.e. convolution, multi-branch [37]) is an additive superposition of pixels, essentially no different from the perceptron in terms of representation space.
@@ 3.2 Dimension of Feature Interaction Space
For example, we consider the Self-Attention in the Transformers [38]. The Self-Attention contains two element-wise multiplications, so it explores a 3-order feature interaction space. This is due to the fact that: (1) in the first stage, in the query-key dot-product attention map computation $Att(\mathbf{x})=Q\cdot K^{T}=W_{Q}\mathbf{x}\cdot\mathbf{x}^{T}W_{K}^{T}$ explores the feature interaction space $\mathcal{A}=\text{span}\left(x_{1}^{2},x_{1}x_{2},\cdots,x_{n-1}x_{n},x_{n}^{2}\right)$. (2) In the second stage, the multiplication between the attention map and the value $y=\sigma(Att(\textbf{x}))\cdot V=\sigma(Att(\textbf{x}))\cdot W_{v}\mathbf{x}$ explores the feature interaction space $\mathcal{S}=\mathcal{A}\otimes\mathbb{R}^{n}=\text{span}\left(x_{1}^{3},x_{1}^{2}x_{2},x_{1}x_{2}x_{3},\cdots,x_{n-1}x_{n}^{2},x_{n}^{3}\right)$, which has $(n+2)(n+1)n/6$ dimensions in line with Eq.(5). Thereby we explain, from the perspective of feature interaction space, why the transformer family of models has, so far, generally outperformed recurrent neural networks and convolutional neural networks in various domains.
@@ 5.1 InfiBlock: Infinite-Dimensional Spatial Feature Interaction
where $K_{rbf}(\vec{Z}^{l}_{a},\vec{Z}^{l}_{b})=exp(-\frac{\|\vec{Z}_{a}^{l}-\vec{Z}_{b}^{l}\|_{2}^{2}}{2})\in\mathbb{R}^{HWC}$, is an RBF kernel with out any hyper-parameter, and $\|\cdot\|_{2}^{2}$ is squared L-2 norm. $\text{Conv}_{a,b}(\hat{X}^{l}W_{a,b})\in\mathbb{R}^{HWC*r}$ expands the number of channels of feature to r times, where $W_{a,b}$ preserve the shape of the feature. The discretization of the feature is performed in a depth-width convolution. $\text{LayerNorm}(\cdot)$ is layer normalization [2] and $\text{MLP}(\cdot)$ is activated by GELU [15].
@@ 6.1 ImageNet Classification
Results. We present our ImageNet experiment results and comparison with baselines in Table 1. Our models achieve competitive performance with state-of-the-art. It is worth noting that InfiNet has about 20% fewer FLOPs [^1] Difference between FLOPs and FLOPS: FLOPs (floating point of operations), the number of floating point operations, is used to measure the complexity of an algorithm/model. This number gets smaller the better. The FLOPs is different from FLOPS (floating point per second), which is a measure of hardware performance. Following other deep-learning architecture works, we use FLOPs to measure the computing demands. Since our model gets a smaller FLOPs, our method is more efficient. than the baseline models with similar parameter scales, but still achieves great performance. It demonstrates the effectiveness of our proposed generation of features in infinite-dimensional interaction spaces. Experiments on isotropic models can be found in Table 3(a).
@@ 7 Conclusion
Limitations
Although we propose the use of kernel methods for infinite-dimensional feature interaction, the only kernel methods we have tried so far are the RBF kernel function and some polynomial kernels of finite dimension. Substitutions utilizing a variety of kernel, including Laplace kernels, exponential kernels, a learnable kernel, etc., can be considered in subsequent studies. Our model has only been performed in some basic computer vision tasks, and validation in language and other modalities still requires some effort. The training of our model is only performed in the supervised learning paradigm, and more training and validation on self-supervised tasks still require effort. In addition, to avoid additional hyperparameter tuning, we fixed the $\sigma$ parameter in the RBF kernel to 1. This may have deprived us of the possibility of exploring the optimal InfiNet, but due to the high cost of training the model, we will leave the impact of this hyperparameter on the InfiNet as a follow-up work.
Broader Social Impact
Acknowledgement
We sincerely thank the reviewers for their valuable suggestions during the review stage.
@@ A.2 More Visualization Comparison
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We attribute the success of modern models to an feature interaction and claim the kernel methods can expand this feature interaction space. And propose the InfiNet. They are included in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: After the Conclusion
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In Section 3 and 4.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In Section 6 and Appendix A.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All datasets we use in this paper are third-party open source.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 6 and Appendix A.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Training for multiple times is too expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Section 6.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This paper is with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: After the Conclusion
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and baselines are cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.